What is the volume (in mL) of 0.250 grams of oxygen gas at a pressure of 0.852 atm and a temperature of -196 °C?

R=0.0821
Answer #1

No. of moles (n) = (0.250 g) / (32 g/mol) = 0.0078 mol

Pressure (P) = 0.852 atm

Temperature (T) =  -196 °C = 77 K

R = 0.0821

PV = nRT

Volume = nRT / P = (0.0078 mol x 0.0821 x 77 K) / (0.852 atm) = 0.0578 L = 57.8 mL

Volume (V) = 57.8 mL